The ratio of the present ages of of Jacob and Nicholas is 2:9. 4 years ago, sum of their ages was 47. how old is Nicholas now?
Answer by ankor@dixie-net.com(22740)

let x = the multiplier
:
The ratio of the present ages of of Jacob and Nicholas is 2:9.
2x = J's age
9x = N's age
:
4 years ago, sum of their ages was 47.
(2x - 4) + (9x - 4) = 47
11x - 8 = 47
11x = 47 + 8
11x = 55
x = 55/11
x = 5 is the multiplier
:
how old is Nicholas now?
N = 9 * 5
N = 45 yrs is N's age